OpenStudy (austinl):

Find the solution of the given initial value problem, and plot its graph. How does the solution behave as \(t\rightarrow\infty\) \(y^{(4)}-4y'''+4y''=0\)

OpenStudy (anonymous):

z10?

OpenStudy (austinl):

My work so far, \(r^4-4r^3+4r^2=0\) \(r^2(r-2)^2=0\) \(r=0,+2,+2\) \(y=c_1+c_2e^{2t}+c_3te^{2t}\) Is this correct?

OpenStudy (anonymous):

i really dont know i am sooo sorry

OpenStudy (anonymous):

theres a medal for my mistake

OpenStudy (austinl):

@ganeshie8 are you any good with these?

OpenStudy (anonymous):

@ganeshie8is very good when it comes to math

OpenStudy (anonymous):

bump the question

OpenStudy (anonymous):

hopefuly it l help u